### v2

# Interpretable Graph Neural Networks for Tabular Data

Abstract
Data in tabular format is frequently occurring in real-world applications. Graph Neural Networks (GNNs) have recently been extended to effectively handle such data, allowing feature interactions to be captured through representation learning. However, these approaches essentially produce black-box models, in the form of deep neural networks, precluding users from following the logic behind the model predictions. We propose an approach, called IGNNet (Interpretable Graph Neural Network for tabular data), which constrains the learning algorithm to produce an interpretable model, where the model shows how the predictions are exactly computed from the original input features. A large-scale empirical investigation is presented, showing that IGNNet is performing on par with state-of-the-art machine-learning algorithms that target tabular data, including XGBoost, Random Forests, and TabNet. At the same time, the results show that the explanations obtained from IGNNet are aligned with the true Shapley values of the features without incurring any additional computational overhead.
1 Introduction
In some application domains, e.g., medicine and law, predictions made by machine learning models need justification for legal and ethical considerations (Lakkaraju et al. 2017; Goodman and Flaxman 2017). In addition, users may put trust in such models only with a proper understanding of the reasoning behind the predictions. A direct solution is to use learning algorithms that produce interpretable models, such as logistic regression (Berkson 1944), which provides both local (instance-specific) and global (model-level) explanations for the predictions. However, such algorithms often result in a substantial loss in predictive performance compared to algorithms that generate black-box models, e.g., XGBoost (Chen and Guestrin 2016), Random Forests (Breiman 2001), and deep learning algorithms (Pintelas, Livieris, and Pintelas 2020; Mori and Uchihira 2019). Post-hoc explanation techniques, e.g., SHAP (Lundberg and Lee 2017), LIME (Ribeiro, Singh, and Guestrin 2016), and Anchors (Ribeiro, Singh, and Guestrin 2018), have been put forward as tools to explain predictions of the black-box models. However, the explanations provided by such techniques are limited in that they either do not show how exactly the predictions are computed, but merely present feature scores, such as LIME and SHAP, or come with no guarantees on the fidelity, i.e., that the provided explanation agrees with the underlying model (Yeh et al. 2019; Delaunay, Galárraga, and Largouët 2020). As extensively argued in (Rudin 2019), there are hence several reasons to consider generating interpretable models in the first place, if trustworthiness is a central concern.
Graph Neural Networks (GNNs) have emerged as a powerful framework for representation learning of graph-structured data (Xu et al. 2019). The application of GNNs has been extended to tabular data, where a GNN can be used to learn an enhanced representation for the data points (rows) or to model the interaction between different features (columns). TabGNN (Guo et al. 2021) is an example of the first approach, where each data point is represented as a node in a graph. In comparison, TabularNet (Du et al. 2021) and Table2Graph (Zhou et al. 2022) follow the second approach, where the first uses a Graph Convolutional Network to model the relationships between features, and the second learns a probability adjacency matrix for a unified graph that models the interaction between features of the data points. GNNs can also be combined with other algorithms suited for tabular data, e.g., as in BGNN (Ivanov and Prokhorenkova 2021), which combines gradient-boosted decision trees and a GNN in one pipeline, where the GNN addresses the graph structure and the gradient-boosted decision trees handle the heterogeneous features of the tabular data. To the best of our knowledge, all previous approaches to using GNNs for tabular data result in black-box models and they are hence associated with the issues discussed above when applied in contexts with strong requirements on trustworthiness. In this work, we propose a novel GNN approach for tabular data, with the aim to eliminate the need to apply post-hoc explanation techniques without sacrificing predictive performance.
The main contributions of this study are:
  • a novel approach, called Interpretable Graph Neural Network for tabular data (IGNNet), that exploits powerful graph neural network models while still being able to show exactly how the prediction is derived from the input features in a transparent way
  • a large-scale empirical investigation evaluating the explanations of IGNNet as well as comparing the predictive performance of IGNNet to state-of-the-art approaches for tabular data; XGBoost, Random Forests, and multi-layer perceptron (MLP), as well as to an algorithm generating interpretable models; TabNet (Arik and Pfister 2021)
  • an ablation study comparing the performance of the proposed approach to a black-box version, i.e., not constraining the learning algorithm to produce transparent models for the predictions
In the next section, we briefly review related work. In Section 3, we describe the proposed interpretable graph neural network. In Section 4, results from a large-scale empirical investigation are presented and discussed, in which the explanations of the proposed method are evaluated and the performance is compared both to interpretable and powerful black-box models. Section 5 discusses the limitations of the proposed approach. Finally, in the concluding remarks section, we summarize the main conclusions and point out directions for future work.
2 Related Work
In this section, we provide some pointers to self-explaining (regular and graph) neural networks, and briefly discuss their relation to model-agnostic explanation techniques and interpretable models. We also provide pointers to work on interpretable deep learning approaches for tabular data.
2.1 Self-Explaining Neural Networks
Several approaches to generating so-called self-explaining neural networks have been introduced in the literature; in addition to generating a prediction model, they all incorporate a component for explaining the predictions. They can be seen as model-specific explanation techniques, in contrast to model-agnostic techniques, such as LIME and SHAP, but sharing the same issues regarding fidelity and lack of detail regarding the exact computation of the predictions.
Approaches in this category include the method in (Lei, Barzilay, and Jaakkola 2016), which is an early self-explaining neural network for text classification, the Contextual Explanation Network (CEN) (Al-Shedivat, Dubey, and Xing 2022), which generates explanations using intermediate graphical models, the Self-explaining Neural Network (SENN) (Alvarez Melis and Jaakkola 2018), which generalizes linear classifiers to neural networks using a concept autoencoder, and the CBM-AUC (Sawada and Nakamura 2022), which improves the efficiency of the former by replacing the decoder with a discriminator. Some approaches generate explanations in the form of counterfactual examples, e.g., CounterNet (Guo, Nguyen, and Yadav 2021), and Variational Counter Net (VCNet) (Guyomard, Fessant, and Guyet 2022). Again, such explanations do not provide detailed information on how the original predictions are computed and how exactly the input features affect the outcome.
2.2 Self-Explaining Graph Neural Networks
The Self-Explaining GNN (SE-GNN) (Dai and Wang 2021) uses similarities between nodes to make predictions on the nodes’ labels and provide explanations using the most similar K nodes with labels. ProtGNN (Zhang et al. 2022) also computes similarities, but between the input graph and prototypical graph patterns that are learned per class. Cui et al. (Cui et al. 2022) proposed a framework to build interpretable GNNs for connectome-based brain disorder analysis that resembles the signal correlation between different brain areas.
Similar to the (standard) self-explaining neural networks, the approaches that target graph neural networks provide abstract views of the predictions; the users cannot trace the exact computations, in contrast to when using interpretable models, which in principle allows for the inferences to be executed by hand.
2.3 Interpretable Deep Learning for Tabular Data
In an endeavor to provide an interpretable regression model for tabular data while retaining the performance of deep learning models, and inspired by generalized linear models (GLM), LocalGLMnet was proposed to make the regression parameters of a GLM feature dependent, allowing for quantifying variable importance and also conducting variable selection (Richman and Wüthrich 2022). TabNet (Arik and Pfister 2021) is another interpretable method proposed for tabular data learning, which employs a sequential attention mechanism and learnable masks for selecting a subset of meaningful features to reason from at each decision step. The feature selection is instance-based, i.e., it differs from one instance to another. The feature selection masks can be visualized to highlight important features and show how they are combined. However, it is not obvious how the features are actually used to form the predictions.
3 The Proposed Approach: IGNNet
This section describes the proposed method to produce an interpretable model using a graph neural network. We first outline the details of a GNN for graph classification and then show how it can be constrained to produce interpretable models. Afterward, we show how it can be applied to tabular data. Finally, we show how the proposed approach can maintain both interpretability and high performance.
3.1 Interpretable Graph Neural Network
The input to a GNN learning algorithm is a set of graphs denoted by $\mathcal{G}=(V,E,X,\mathcal{W})$, consisting of a set of nodes $V$, a set of edges $E$, a set of node feature vectors $X$, and a set of edge weights $\mathcal{W}$, where $V=\{v_{1},\ldots,v_{N}\}$, $E\subseteq\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$, $X=\{\textbf{x}_{1},\ldots,\textbf{x}_{N}\}$, and the weight of edge $(v_{i},v_{j})$ is represented by a scalar value $\delta_{i,j}$ in the set of edge weights $\mathcal{W}$, where $\delta_{i,j}=\mathcal{W}(i,j)$. A GNN algorithm learns a representation vector $\textbf{h}_{i}$ for each node $v_{i}$, which is initialized as $\textbf{h}_{i}^{(0)}=\textbf{x}_{i}$. The key steps in a GNN for graph classification can be summarized by the following two phases (Xu et al. 2019):
  1. (a)
    Message Passing: Each node passes a message to the neighboring nodes, then aggregates the passed information from the neighbors. Finally, the node representation is updated with the aggregated information. A neural network can also be used to learn some message functions between nodes. The message passing phase can be formulated as:
    $\textbf{h}_{i}^{(l+1)}=\varphi\left(\textbf{w}^{(l+1)}\left(\delta_{i,i}\textbf{h}_{i}^{(l)}+\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}\right)\right)$ (1)
    where $\delta_{i,u}$ is the weight assigned to the edge between node $v_{i}$ and node $v_{u}$. $\textbf{h}_{i}^{(l)}$ is the hidden representation of the node $v_{i}$ in the $l$-th layer, $\textbf{w}^{(l+1)}$ represents the learnable parameters, and $\varphi$ is a non-linearity function.
    The adjacency matrix A of size $|V|\times|V|$ contains the edge weights and can be normalized similar to a Graph Convolutional Network (GCN) (Kipf and Welling 2017) as shown in (2).
    $\tilde{\textit{{A}}}=\textit{{D}}^{-\frac{1}{2}}\textit{{A}}\textit{{D}}^{-\frac{1}{2}}$ (2)
    Here D is the degree matrix $\textit{{D}}_{ii}=\sum_{j}\textit{{A}}_{ij}$ (Kipf and Welling 2017).
  2. (b)
    Graph Pooling (Readout): A representation of the whole graph $\mathcal{G}$ is learned using a simple or advanced function (Xu et al. 2019), e.g, sum, mean, or MLP.
The whole graph representation obtained from the graph pooling phase can be submitted to a classifier to predict the class of the graph, which can be trained in an end-to-end architecture (Zhang et al. 2018; Ying et al. 2018).
The pooling function can be designed to provide an interpretable graph classification layer. Thus, the final hidden representation of each node is mapped using an injective readout function to a single value, for instance, through a neural network layer or dot product ($\mathcal{R}(\textbf{h}_{i}^{(l+1)}\in\mathbb{R}^{n})=h_{i}\in\mathbb{R}^{1}$), and concatenated to obtain the final representation g of the graph $\mathcal{G}$ where a scalar value in g corresponds to a node in the graph. Consequently, if a set of weights is applied to classify the graph, we can trace the contribution of each node to the predicted outcome, i.e., the user can find out which nodes contributed to the predicted class. For example, g can be used directly as follows:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{\textbf{w}_{i}\textbf{g}_{i}}\right)$ (3)
where $\textbf{w}_{i}$ is the weight assigned to node $v_{i}$ represented in $\textbf{g}_{i}$.The link function is applied to accommodate a valid range of outputs, e.g., the sigmoid function for binary and softmax for multi-class classification. This is equivalent to:
$\hat{y}=\text{link}\left(\sum_{i=1}^{n}{\textbf{w}_{i}\mathcal{R}(\textbf{h}_{i}^{(l+1)})}\right)$ (4)
In the case of binary classification, one vector of weights (w) is applied, and for multiple classes, each class has a separate vector of weights.
3.2 Representing Tabular Data Points as Graphs
Figure 1: An overview of our proposed approach. Each data instance is represented as a graph by embedding the feature values into a higher dimensionality, and the edge between two features (nodes) is the correlation value. Multiple iterations of message passing are then applied. Finally, the learned node representation is projected into a single value, and a whole graph representation is obtained by concatenating the projected values.
The proposed readout function in the previous subsection allows for determining the contribution of each node in a prediction, if a white-box classification layer is used for the latter. Therefore, we propose representing each data instance as a graph where the features are the nodes of that graph and the linear correlation between features are the edge weights, as we assume that not all features are completely independent. The initial representation of a node is a vector of one dimension, and the value is just the feature value, which can be embedded into a higher dimensionality. The idea is outlined in Algorithm 1 and illustrated in Figure 1.
Data: a set of graphs $\mathbb{G}$ and labels $y$
Result: Model parameters $\theta$
Initialize $\theta$
for number of training iterations do
       $\mathcal{L}\leftarrow 0$
       for each $\mathcal{G}_{j}\in\mathbb{G}$ do
             $\textit{{H}}^{(0)}_{j}\leftarrow\mathcal{G}_{j}$
             for each layer l $\in$ messagePassing layers do
                   $\textit{{H}}^{(l+1)}_{j}\leftarrow$messagePassing$(\textit{{H}}^{(l)}_{j})$
             end for
            $\textbf{g}_{j}\leftarrow$readout$(\textit{{H}}^{(l+1)}_{j})$
             $\hat{y}_{j}\leftarrow$predict$(\textbf{g}_{j})$
             $\mathcal{L}\leftarrow\mathcal{L}+$loss$(\hat{y}_{j},y_{j})$
       end for
      Compute gradients $\nabla_{\theta}\mathcal{L}$
       Update $\theta\leftarrow\theta-\nabla_{\theta}\mathcal{L}$
end for
Algorithm 1 IGNNet
In order to make a prediction for a test instance, the data point has to be converted to a graph using the same procedure for building the input graphs to IGNNet, and for which graph node representations are obtained using a GNN with parameters $\theta$. Finally, the output layer is used to form the prediction.
3.3 How can IGNNet achieve high performance while maintaining interpretability?
An expressive GNN can potentially capture complex patterns and dependencies in the graph, allowing nodes to be mapped to distinct representations based on their characteristics and relationships (Li and Leskovec 2022). Moreover, a GNN with an injective aggregation scheme can not only distinguish different structures but also map similar structures to similar representations (Xu et al. 2019). Therefore, if the tabular data are properly presented as graphs, GNNs with the aforementioned expressive capacities can model relationships and interactions between features, and consequently approximate complex non-linear mappings from inputs to predictions. On top of that, it has been shown by (ENNADIR et al. 2023) that GCNs based on 1-Lipschitz continuous activation functions can be improved in stability and robustness with Lipschitz normalization and continuity analysis; similar findings have also been demonstrated on graph attention networks (GAT) (Dasoulas, Scaman, and Virmaux 2021). This property is of particular importance when the application domain endures adversarial attacks or incomplete tabular data.
The proposed readout function in subsection 3.1 can produce an interpretable output layer. However, it does not guarantee the interpretability of the whole GNN without message-passing layers that consistently maintain relevant representations of the input features. Accordingly, we constrain the message-passing layer to produce interpretable models using the following conditions:
  1. 1.
    Each feature is represented in a distinct node throughout the consecutive layers.
  2. 2.
    Each node is bounded to interact with a particular neighborhood, where it maintains correlations with the nodes within that neighborhood. As a result, the aggregated messages could potentially hold significance to the input feature values.
Since the edge weights are the correlation values, they determine the strength and the sign of the messages obtained from a neighborhood, allowing each node to store information not only concerning the original input feature but also features that are correlated with it. The proposed graph pooling function, combined with the constrained message-passing layers that keep representative information about the input features, allows tracking each feature’s contribution at the output layer and also through the message-passing layers all the way to the input features.
4 Empirical Investigation
This section evaluates both the explanations and predictive performance of IGNNet. We begin by outlining the experimental setup, then by evaluating the explanations produced by IGNNet, and lastly, we benchmark the predictive performance.
4.1 Experimental Setup
A GNN consists of one or more message-passing layers, and each layer can have a different design, e.g., different activation functions and batch normalization, which is the intra-layer design level (You, Ying, and Leskovec 2020). There is also the inter-layer design level which involves, for instance, how the message-passing layers are organized into a neural network and if the skip connections are added between layers (You, Ying, and Leskovec 2020). While the intra-layer and inter-layer designs can vary based on the nature of the prediction task, we propose a general architecture for our empirical investigation. [^1] The source code is available here: https://github.com/amrmalkhatib/ignnet However, it is up to the user to modify the architecture of the GNN. The number of message-passing layers, number of units in linear transformations, and other hyperparameters were found based on a quasi-random search and evaluation on development sets of the following three datasets: Churn, Electricity, and Higgs. We have six message-passing layers in the proposed architecture, each with a Relu activation function. Multiple learnable weights are also applied to the nodes’ representation, followed by a Relu function. Besides three batch normalization layers, four skip connections are added as illustrated in Figure 2. After all the GNN layers, we use a feedforward neural network (FNN) to map the multidimensional representation of each node into a single value. In the FNN, we do not include any activation functions in order to keep the mapping linear, but a sigmoid function is applied after the final layer to obtain a value between 0 and 1 for each node. The FNN is composed of 8 layers with the following numbers of units (128, 64, 32, 16, 8, 4, 2, 1) and 3 batch normalization layers after the second, fourth, and sixth hidden layers. After the FNN, the nodes’ final values are concatenated to form a representation of the whole graph (data instance). Finally, the weights that are output are used to make predictions. The GNN is trained end-to-end, starting from the embeddings layer and ending with the class prediction.
Figure 2: IGNNet default architecture. It starts with the embedding layer, a linear transformation from one dimension to 64 dimensions. A Relu activation function follows each message-passing layer and each green block as well. The feedforward network at the end has no activation functions between layers to ensure a linear transformation into a single value. A sigmoid activation function follows the feedforward network to obtain the final value for each feature between 0 and 1.
We also provide an opaque variant of IGNNet (OGNNet, opaque graph neural net for tabular data), where the FNN, along with the output layer, is replaced by an MLP of one hidden layer with 1024 hidden units and a Relu activation function. All the learned node representations just before the FNN are concatenated and passed to the MLP for class prediction. The OGNNet is introduced to determine, by an ablation study, how much predictive performance we may lose by squashing the learned multidimensional representation of the nodes into scalar values and applying a white box classifier instead of a black box.
In the experiments, 35 publicly available datasets are used. [^2] All the datasets were obtained from https://www.openml.org Each dataset is split into training, development, and test sets. The development set is used for overfitting detection and early stopping of the training process, the training set is used to train the model, and the test set is used to evaluate the model. [^3] Detailed information about each dataset is provided in the appendix. For a fair comparison, all the compared learning algorithms are trained without hyperparameters tuning using the default settings on each dataset. In cases where the learning algorithm does not employ the development set to measure performance progress for early stopping, the development and training subsets are combined into a joint training set. The adjacency matrix uses the correlation values computed on the training data split. The weight on edge from the node to itself (self-loop) is a user-adjustable hyperparameter, constitutes between 70% to 90% (on average) of the weighted summation to keep a strong message per node that does not fade out with multiple layers of message-passing. Weak correlation values are excluded from the graph, so if the absolute correlation value is below 0.2, the edge is removed unless no correlation values are above 0.2; in case of the latter, the procedure is repeated using a reduced threshold of 0.05.444In the technical appendix, we provide an ablation study on the effects of different hyperparameter preferences on predictive performance. The Pearson correlation coefficient (Pearson 1895) is used to estimate the linear relationship between features. In the data preprocessing step, the categorical features are binarized using one-hot encoding, and all the feature values are normalized using min-max normalization (the max and min values are computed on the training split). The normalization keeps the feature values between 0 and 1, which is essential for the IGNNet to have one scale for all nodes.
The following algorithms are also evaluated in the experiments: XGBoost, Random Forests, MLP and TabNet. XGBoost and Random Forests are trained on the combined training and development sets. The MLP has two layers of 1024 units with Relu activation function and is trained on the combined training and development sets with early stopping and 0.1 validation fraction. TabNet is trained with early stopping after 20 consecutive epochs without improvement on the development set, and the best model is used in the evaluation.
For imbalanced binary classification datasets, we oversample the minority class in the training set to align the size with the majority class. All the compared algorithms are trained using the oversampled training data. While for multi-class datasets, no oversampling is conducted. The area under the ROC curve (AUC) is used to measure the predictive performance, as it is not affected by the decision threshold. For the multi-class datasets, weighted AUC is calculated, i.e., the AUC is computed for each class against the rest and weighted by the support.
Figure 3: Comparison of KernelSHAP’s approximations and the importance scores obtained from IGNNet. We measure the similarity of KernelSHAP’s approximations to the scores of IGNNet at each iteration of data sampling and evaluation of KernelSHAP. KernelSHAP exhibits improvement in approximating the scores derived from IGNNet with more data sampling.
4.2 Evaluation of Explanations
The feature scores produced by IGNNet should ideally reflect the contribution of each feature toward the predicted outcome and, therefore, they should be equivalent to the true Shapley values. As it has been shown that KernelSHAP converges to the true Shapley values when provided with an infinite number of samples (Covert and Lee 2021; Jethani et al. 2022), it is anticipated that the explanations generated by KernelSHAP will progressively converge to more similar values to the scores of IGNNet as the sampling process continues. This convergence arises from KernelSHAP moving towards the true values, while the scores of IGNNet are expected to align with these true values. To examine this conjecture, we explain IGNNet using KernelSHAP and measure the similarity between KernelSHAP’s explanations and IGNNet’s scores following each iteration of data sampling and KernelSHAP evaluation. [^5] KernelSHAP experiments were conducted using the following open-source implementation: https://github.com/iancovert/shapley-regression For the feasibility of the experiment, 500 examples are randomly selected from the test set of each dataset to be explained. The cosine similarity and Spearman rank-order correlation are used to quantify the similarity between explanations. The cosine similarity measures the similarity in the orientation (Han, Kamber, and Pei 2012), while the Spearman rank-order measures the similarity in ranking the importance scores (Rahnama et al. 2021).
The results demonstrate a general trend wherein KernelSHAP’s explanations converge to more similar values to IGNNet’s scores across various data instances and the 35 datasets, as depicted in Figure 3. The consistent convergence to more similar values clearly indicates that IGNNet provides transparent models with feature scores aligned with the true Shapley values. [^6] The complete results of the 35 datasets are provided in the technical appendix.  [^7] We provide illustrations of individual explanations in the technical appendix.
4.3 Evaluation of Predictive Performance
Detailed results for IGNNet and the five competing algorithms on the 35 datasets are shown in Table 1. The ranking of the five algorithms across 35 datasets, based on their AUC values, reveals OGNNet to exhibit superior performance, claiming the top position, closely followed by IGNNet and XGBoost. In order to investigate whether the observed differences are statistically significant, the Friedman test (Friedman 1939) was employed, which indeed allowed to reject the null hypothesis, i.e., that there is no difference in predictive performance, as measured by the AUC, at the 0.05 level. The result of subsequently applying the post-hoc Nemenyi test (Nemenyi 1963) to determine what pairwise differences are significant, again at the 0.05 level, is summarized in Figure 4. However, the result shows no specific significant pairwise differences between any of the compared algorithms. Furthermore, the results show that using IGNNet instead of the black-box variant, OGNNet, does not significantly reduce the predictive performance while maintaining performance at the level of other powerful algorithms for tabular data, e.g., XGBoost and Random Forests.
Figure 4: The average rank of the compared classifiers on the 35 datasets with respect to the AUC (a lower rank is better), where the critical difference (CD) represents the largest difference that is not statistically significant.
Table 1: The AUC of IGNNet, OGNNet, MLP, Random Forests, and XGBoost. The best-performing model is colored in blue, and the second best-performing is colored in light blue.
Dataset IGNNet OGNNet TabNet MLP Random Forests XGBoost Abalone 0.881 0.883 0.857 0.877 0.876 0.869 Ada Prior 0.905 0.888 0.848 0.877 0.885 0.894 Adult 0.917 0.915 0.919 0.881 0.907 0.931 Bank 32 nh 0.887 0.887 0.881 0.859 0.876 0.874 Covertype 0.984 0.988 0.969 0.861 0.995 0.967 Credit Card Fraud 0.987 0.966 0.969 0.913 0.914 0.975 Delta Ailerons 0.977 0.977 0.974 0.977 0.978 0.977 Electricity 0.901 0.929 0.894 0.928 0.97 0.973 Elevators 0.951 0.948 0.95 0.95 0.913 0.943 HPC Job Scheduling 0.908 0.921 0.775 0.908 0.952 0.955 Fars 0.956 0.958 0.954 0.955 0.949 0.962 1st Order Theorem Proving 0.776 0.808 0.495 0.805 0.854 0.858 Helena 0.875 0.889 0.884 0.897 0.855 0.875 Heloc 0.783 0.787 0.772 0.783 0.778 0.775 Higgs 0.762 0.785 0.804 0.774 0.793 0.797 Indian Pines 0.984 0.992 0.99 0.973 0.979 0.987 Jannis 0.856 0.857 0.867 0.861 0.861 0.872 JM1 0.739 0.725 0.711 0.728 0.747 0.733 LHC Identify Jets 0.941 0.941 0.944 0.861 0.935 0.941 Madelon 0.906 0.718 0.501 0.668 0.79 0.891 Magic Telescope 0.907 0.921 0.927 0.929 0.934 0.928 MC1 0.957 0.904 0.89 0.853 0.844 0.943 Mozilla4 0.954 0.961 0.971 0.963 0.988 0.99 Microaggregation2 0.778 0.792 0.752 0.782 0.768 0.781 Numerai28.6 0.526 0.534 0.52 0.518 0.519 0.514 Otto Group Product 0.96 0.971 0.968 0.972 0.973 0.974 PC2 0.881 0.815 0.844 0.571 0.55 0.739 Phonemes 0.922 0.96 0.898 0.93 0.964 0.953 Pollen 0.492 0.508 0.509 0.496 0.489 0.467 Satellite 0.998 0.993 0.911 0.992 0.998 0.992 Scene 0.994 0.989 0.986 0.992 0.983 0.982 Speed Dating 0.853 0.835 0.797 0.822 0.845 0.86 Telco Customer Churn 0.858 0.845 0.841 0.783 0.84 0.843 Vehicle sensIT 0.918 0.918 0.917 0.914 0.912 0.916 Waveform-5000 0.965 0.962 0.933 0.965 0.959 0.957
4.4 Computational Cost
The computational cost varies based on the architecture of the graph neural network, which can be altered and determined based on the predictive task and the dataset. The cost relative to the conventional graph neural networks remains the same, i.e., the cost does not increase with the proposed approach. The computational cost also depends on the number of features in the dataset. Therefore, the user can decide on the suitable architecture and the acceptable computational cost. Detailed information about the inference time is available in the appendix.
5 Limitations
The proposed approach targets tabular data primarily consisting of numerical features. However, datasets with substantial categorical features impose two challenges. Firstly, Pearson correlation is inadequate in handling correlation between categorical and numerical features or between two categorical features. Secondly, we use one-hot encoding for categorical features, which can result in an exponential growth in dimensionality when dealing with nominal features with numerous categories. Another limitation arises when input features lack correlation, resulting in a completely disconnected graph.
6 Concluding Remarks
We have proposed IGNNet, an algorithm for tabular data classification, which exploits graph neural networks to produce transparent models. In contrast to post-hoc explanation techniques, IGNNet does not approximate or require costly computations, but provides the explanation while computing the prediction, and where the explanation prescribes exactly how the prediction is computed.
We have presented results from a large-scale empirical investigation, in which IGNNet was evaluated with respect to explainability and predictive performance. IGNNet was shown to generate explanations with feature scores aligned with the Shapley values without further computational cost. IGNNet was also shown to achieve a similar predictive performance as XGBoost, Random Forests, TabNet, and MLP, which are all well-known for their ability to generate high-performing models.
One direction for future research is to explore approaches to model feature interaction in the adjacency matrix that go beyond linear correlations. Understanding how such non-linear interactions between features may impact the model’s interpretability could be an intriguing area of exploration. A second direction is to investigate alternative encoders for categorical features rather than relying on one-hot encoding. It would also be interesting to extend IGNNet to handle non-tabular datasets, including images and text, which would require entirely different approaches to representing each data point as a graph. Another important direction for future work is to use IGNNet for studying possible adversarial attacks on a predictive model. Finally, an important direction would be to complement the empirical evaluation with user-grounded evaluations, e.g., measuring to what extent certain tasks could be more effectively and efficiently solved when the users are provided with a transparent model that shows how the prediction has been computed from the input.
7 Acknowledgement
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Appendix
Appendix B Illustration of Explanations
In this section, we show how the computed feature scores by IGNNet can be used to understand the feature contributions toward a specific prediction. Note that this is done in exactly the same way as how one would interpret the predictions of a logistic regression model or the feature importance scores generated by the SHAP explainer (Lundberg and Lee 2017). As the computed feature scores reveal exactly how IGNNet formed the prediction, the user can directly see which features have the greatest impact on the final prediction, and possibly also how they may be modified to affect the outcome. To demonstrate this, we present the feature scores for predictions made by IGNNet using two examples from the Adult dataset and the Churn dataset. In the following illustrations, we display the feature scores centered around the bias value, which, when summed with the bias, will produce the exact outcome of IGNNet if the sigmoid function is applied. The scores are sorted according to their absolute values, and only the top 10 features are plotted for ease of presentation. A displayed score $\tau_{i}$ of feature $x_{i}$ represents all the weights and the computations applied to the input value, as shown in equation 5.
$\tau_{i}=$ $\textbf{w}_{i}f(\varphi(\textbf{w}^{(l+1)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l)}$ (5)
$+\delta_{i,i}\varphi(\textbf{w}^{(l)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{h}_{u}^{(l-1)}$
$+\delta_{i,i}(...\varphi(\textbf{w}^{(1)}(\sum_{u\in\mathcal{N}(i)}\delta_{i,u}\textbf{x}_{u}+\delta_{i,i}\textbf{x}_{i})...))))))$
Figure 5: Explanation to a single prediction on Adult dataset.
Figure 6: The explanation of a single prediction on Churn dataset.
Figure 7: The correlation matrix of the features of the Churn dataset.
The first example, from the Adult dataset, shown in Figure 5(a). IGNNet predicted the negative class ($\leq$ 50K) with a narrow margin (0.495). The explanation shows that a single feature (capital-gain=2885) has the highest contribution compared to any other feature value. In the training data, the capital-gain has a maximum value of 99999.0, a minimum value of 0, a mean value of 1068.36, and a 7423.08 standard deviation. To test if the explanation reflects the actual reasoning of IGNNet, we raise the capital-gain value by a smaller value than the standard deviation to be 6000 while leaving the remaining feature values constant, and it turns out to be enough to alter the prediction to a positive ($>$ 50K) with 0.944 as the predicted value. We can also see that the negative score of the capital-gain feature went from -3.67 in the original instance (5(a)) to -0.82 in the modified instance, as shown in Figure 5(b). So the user can adjust the value of an important feature as much as needed to alter the prediction.
The data point, from the Churn dataset, has a positive prediction with a narrow margin (0.565). Consequently, the reduction of the top positively important feature (total day charge), as illustrated in Fig. 6(a), may be enough to obtain a negative prediction. The total day charge has a maximum value of 59.76, a minimum of 0.44, a mean of 30.64, and a 9.17 standard deviation in the training set. However, the total day charge is highly correlated with the total day minutes by more than 0.99, as shown in Fig. 7. The high correlation effect is obvious in the outcome when the total day charge value is reduced by one standard deviation, from 44.56 to 35.39, as the scores of both total day charge and total day minutes drop from 1.06 and 0.99 to 0.72 and 0.5, respectively, as shown in Fig. 6(b). Moreover, the sum of feature scores and bias falls below 0.5 after the sigmoid function, resulting in a negative prediction with a predicted value of 0.297.
Appendix C Ablation Study
In this section, we conduct three sets of experiments to determine the effects of different hyperparameter choices on predictive performance. In the first set, we study the effects of changing the weight on the edge from a feature to itself in the adjacency matrix (self-loop). In the second set, we inspect the effect of enforcing a threshold on the correlation value between two features to be included as an edge weight. Finally, the third set of experiments determines the effect of using a different number of message-passing layers.
Weight on Self-Loop
In the experimental setup, we evaluated IGNNet using a value that forms 70% up to 90% of the weighted summation on the self-loop. In this set of experiments, we evaluate IGNNet using self-loops that form more than 90% of the weighted summation, then we test with a fixed value of 1 for the self-loop, and finally, we evaluate with 0 self-loops. The detailed results are available in table 2. The null hypothesis that there is no difference in the predictive performance between IGNNet in the default settings, and IGNNet with the three new variations of the self-loops, has been tested using the Friedman test followed by the pairwise Nemenyi tests. The null hypothesis can be rejected at 0.05 level for the difference between the default settings and the tested variations of self-loops, as the default settings showed significantly better performance. The result of the post hoc tests are summarized in Fig. 8.
Figure 8: The average rank of IGNNet in the default settings and other IGNNet models with different values for the self-loops on the 30 datasets. The ranking has been done for the predictive performance measured in the AUC.
Threshold on Correlation Values
In the experimental setup, we set a threshold for a correlation value to be added as an edge between two features, which is 0.2 in general and 0.05 in the case of small correlation values. In this set of experiments, we compare the performance of IGNNet with the abovementioned thresholds and without any thresholds. To test the null hypothesis that there is no difference in the predictive performance, as measured by AUC, between the two variations, and since we compare two models, the Wilcoxon signed-rank test (Wilcoxon 1945) is used here. The null hypothesis may be rejected at the 0.05 level, showing that setting a threshold for the edge weight helps with improving the predictive performance of IGNNet. The results are available in Table 2.
The Number of Message-Passing Layers
In this final set of experiments, we investigate the effect of reducing the number of the message-passing layers and the total number of parameters in the IGNNet on the predictive performance. We compare IGNNet (with the default architecture) to a reduced version of three message-passing layers and another version with just one message-passing layer. For the three layers version, we use the architecture of IGNNet as mentioned in the experimental setup, but we keep the first three message-passing layers up to the first batch normalization, then we remove the following three message-passing layers and their linear transformations before the final fully-connected feedforward network. The model with one message-passing layer has only the first layer with the subsequent transformations, then all the subsequent layers are removed up until the fully-connected feedforward network is.
Similar to what has been done in the previous subsections, we test the null hypothesis of no difference in the predictive performance between the three architectures using the Friedman test, followed by pairwise post hoc Nemenyi tests. The first test rejects the null hypothesis, and the result of the Nemenyi tests are outlined in Fig. 9. The results in Table 2 demonstrate that the predictive performance generally improves with more layers of message-passing and linear transformations.
Figure 9: The average rank of IGNNet with 6, 3, and 1 message-passing layers on the 30 datasets. The ranking has been done for the predictive performance measured in the AUC.
Table 2: The results of the ablation study that compares the predictive performance of IGNNet in the default settings to IGNNet with different weights on the self-loop, IGNNet without a threshold on the correlation value to be included as an edge in the adjacency matrix, and IGNNet with different numbers of message-passing layers. The best-performing model is colored in blue, and the second best-performing is colored in light blue.
Self-Loop Experiments Number of Layers Experiments Dataset Default Settings Above 90% Self-Loop Value = 1 Self-Loop Value = 0 Without Correlation Threshold 3 Layers 1 Layer Abalone 0.88 0.88 0.844 0.842 0.88 0.874 0.848 Ada Prior 0.905 0.905 0.891 0.88 0.896 0.904 0.903 Adult 0.917 0.914 0.916 0.88 0.916 0.917 0.911 Bank 32 nh 0.887 0.889 0.854 0.815 0.886 0.877 0.881 Breast Cancer 0.999 0.999 0.998 0.998 0.999 0.998 0.998 Churn 0.88 0.865 0.871 0.9 0.905 0.872 0.868 Credit Card Fraud 0.987 0.974 0.942 0.913 0.952 0.938 0.937 Delta Ailerons 0.977 0.977 0.977 0.976 0.977 0.976 0.977 Delta Elevators 0.951 0.952 0.946 0.946 0.951 0.95 0.949 Electricity 0.898 0.871 0.881 0.844 0.866 0.845 0.838 Elevators 0.951 0.946 0.923 0.928 0.931 0.896 0.847 Higgs 0.762 0.791 0.696 0.707 0.758 0.727 0.717 JM1 0.739 0.729 0.702 0.71 0.734 0.721 0.723 Madelon 0.906 0.717 0.91 0.895 0.855 0.807 0.686 Magic Telescope 0.907 0.929 0.89 0.91 0.919 0.901 0.873 Mozilla4 0.954 0.953 0.954 0.965 0.963 0.942 0.862 MC1 0.957 0.909 0.953 0.935 0.933 0.904 0.961 Numerai28.6 0.526 0.52 0.52 0.521 0.52 0.525 0.525 PC2 0.881 0.883 0.865 0.859 0.827 0.849 0.875 Phishing 0.99 0.987 0.989 0.993 0.992 0.988 0.987 Phonemes 0.922 0.919 0.924 0.89 0.941 0.915 0.895 Pollen 0.525 0.491 0.516 0.524 0.483 0.489 0.498 Satellite 0.998 0.996 0.928 0.931 0.984 0.996 0.995 Scene 0.994 0.959 0.969 0.932 0.992 0.957 0.941 Spambase 0.984 0.98 0.974 0.963 0.983 0.985 0.982 Speed Dating 0.853 0.844 0.825 0.812 0.824 0.82 0.814 Telco Customer Churn 0.858 0.86 0.841 0.844 0.858 0.852 0.853 Tic Tac Toe 0.846 0.828 0.841 0.864 0.876 0.831 0.803 Vehicle sensIT 0.918 0.916 0.915 0.91 0.915 0.915 0.914 Waveform-5000 0.965 0.962 0.966 0.964 0.961 0.96 0.96
Appendix D Transparency Evaluation
In this section, we demonstrate the detailed results of the explanations evaluation experiment using 35 datasets. The results show a general trend across the 35 datasets where the explanations obtained using KernelSHAP converge to more similar values to the feature scores produced by IGNNet, given more sampled data. The results are displayed in Figure 10, Figure 11, and Figure 12.
Figure 10: Comparison of KernelSHAP’s approximations and the importance scores obtained from IGNNet. We measure the similarity of KernelSHAP’s approximations to the scores of IGNNet at each iteration of data sampling and evaluation of KernelSHAP. KernelSHAP exhibits improvement in approximating the scores derived from IGNNet with more data sampling.
Figure 11: Comparison of KernelSHAP’s approximations and the importance scores obtained from IGNNet. We measure the similarity of KernelSHAP’s approximations to the scores of IGNNet at each iteration of data sampling and evaluation of KernelSHAP. KernelSHAP exhibits improvement in approximating the scores derived from IGNNet with more data sampling.
Figure 12: Comparison of KernelSHAP’s approximations and the importance scores obtained from IGNNet. We measure the similarity of KernelSHAP’s approximations to the scores of IGNNet at each iteration of data sampling and evaluation of KernelSHAP. KernelSHAP exhibits improvement in approximating the scores derived from IGNNet with more data sampling.
Appendix E Information about the Used Datasets and Specifications of the Hardware
This subsection provides a summary of the datasets utilized in the experiments. In Table 3, we provide information about the used datasets, including the number of classes (Num of Classes), the number of features (Num of Features), the size of the dataset, the size of the training, validation, and test splits, as well as, the used correlation threshold of each dataset, the weight on the self-loop (Self-Loop Wt.) and finally the ID of each dataset on OpenML.
The experiments have been performed in a Python environment on an Intel(R) Core(TM) i9-10885H CPU @ 2.40GHz system with 64.0 GB of RAM, and the GPUs are NVIDIA® GeForce® GTX 1650 Ti with 4 GB GDDR6, and NVIDIA® GeForce® GTX 1080 Ti with 8 GB. All the software and package dependencies are documented with the source code.
Table 3: The dataset information.
Dataset Classes Features Dataset Size Training Set Dev. Set Test Set Corr. Threshold Self-Loop Wt. Training Epochs OpenML ID Abalone 2 8 4177 2506 836 835 0.2 20 220 720 Ada Prior 2 14 4562 2737 913 912 0.2 4 216 1037 Adult 2 14 48842 43957 2443 2442 0.2 4 297 1590 Bank 32 nh 2 32 8192 5734 1229 1229 0.2 2 540 833 Covertype 7 54 581012 524362 27599 29051 0.2 10 300 1596 Credit Card Fraud 2 30 284807 270566 7121 7120 0.2 3 29 42175 Delta Ailerons 2 5 7129 3564 1783 1782 0.2 2 380 803 Electricity 2 8 45312 36249 4532 4531 0.2 3 396 151 Elevators 2 18 16599 11619 2490 2490 0.2 4 353 846 HPC Job Scheduling 4 7 4331 2598 867 866 0.05 10 435 43925 Fars 8 29 100968 80774 10097 10097 0.05 10 301 40672 1st Order Theorem Prov. 6 51 6118 3915 979 1224 0.2 30 848 1475 Helena 100 27 65196 41724 10432 13040 0.2 10 550 41169 Heloc 2 22 10000 7500 1250 1250 0.2 20 234 45023 Higgs 2 28 98050 88245 4903 4902 0.05 4 394 23512 Indian Pines 8 220 9144 5852 1463 1829 0.2 400 394 41972 Jannis 4 54 83733 53588 13398 16747 0.05 20 300 41168 JM1 2 21 10885 8708 1089 1088 0.2 50 187 1053 LHC Identify Jets 5 16 830000 749075 39425 41500 0.2 10 300 42468 Madelon 2 500 2600 1560 520 520 0.05 4 199 1485 Magic Telescope 2 10 19020 15216 1902 1902 0.2 4 397 1120 MC1 2 38 9466 7478 994 994 0.2 80 198 1056 Mozilla4 2 5 15545 12436 1555 1554 0.2 2 280 1046 Microaggregation2 5 20 20000 12800 3200 4000 0.2 15 599 41671 Numerai28.6 2 21 96320 86688 4816 4816 0.2 20 36 23517 Otto Group Product 9 93 61878 39601 9901 12376 0.05 30 270 45548 PC2 2 36 5589 3353 1118 1118 0.2 60 37 1069 Phonemes 2 5 5404 3782 811 811 0.2 1 800 1489 Pollen 2 5 3848 2308 770 770 0.2 4 351 871 Satellite 2 36 5100 2805 1148 1147 0.2 60 287 40900 Scene 2 299 2407 1203 602 602 0.2 40 86 312 Speed Dating 2 120 8378 5864 1257 1257 0.2 10 58 40536 Telco Customer Churn 2 19 7043 4930 1057 1056 0.2 3 116 42178 Vehicle sensIT 2 100 98528 88675 4927 4926 0.2 10 172 357 waveform-5000 2 40 5000 3000 1000 1000 0.2 4 97 979
Appendix F Inference Time
The inference time has been reported using an Intel(R) Core(TM) i9-10885H CPU @ 2.40GHz system with 64.0 GB of RAM, and the GPUs are NVIDIA® GeForce® GTX 1650 Ti with 4 GB GDDR6. The time is reported as the average of 100 different data objects in milliseconds. The results are shown in Table 4.
Table 4: The dataset information.
Dataset Classes Features OHE Nodes Non-Zero Edges Inference Time Abalone 2 8 10 100 3.74 $\pm$ 0.77 Ada Prior 2 14 100 284 4.17 $\pm$ 0.7 Adult 2 14 105 283 4.11 $\pm$ 0.69 Bank 32 nh 2 32 32 38 4.06 $\pm$ 0.72 Covertype 7 54 98 488 9.03 $\pm$ 1.35 Credit Card Fraud 2 30 30 50 3.93 $\pm$ 0.68 Delta Ailerons 2 5 5 11 3.91 $\pm$ 0.67 Electricity 2 8 14 38 3.87 $\pm$ 0.73 Elevators 2 18 18 112 4.31 $\pm$ 0.55 HPC Job Scheduling 4 7 26 258 8.48 $\pm$ 1.81 Fars 8 29 107 2295 8.18 $\pm$ 1.11 1st Order Theorem Prov. 6 51 51 1251 8.34 $\pm$ 0.92 Helena 100 27 27 293 4.61 $\pm$ 0.57 Heloc 2 22 22 370 5.48 $\pm$ 0.51 Higgs 2 28 28 244 6.58 $\pm$ 0.6 Indian Pines 8 220 220 40948 10.7 $\pm$ 1.53 Jannis 4 54 54 542 4.15 $\pm$ 0.7 JM1 2 21 21 415 9.92 $\pm$ 0.79 LHC Identify Jets 5 16 16 184 3.85 $\pm$ 0.76 Madelon 2 500 500 12566 10.42 $\pm$ 2.02 Magic Telescope 2 10 10 54 5.88 $\pm$ 0.53 MC1 2 38 38 1142 4.39 $\pm$ 0.71 Mozilla4 2 5 5 7 4.03 $\pm$ 0.55 Microaggregation2 5 20 20 138 3.82 $\pm$ 0.7 Numerai28.6 2 21 21 231 4.68 $\pm$ 0.618 Otto Group Product 9 93 93 3581 4.06 $\pm$ 0.82 PC2 2 36 36 1040 4.37 $\pm$ 0.73 Phonemes 2 5 5 9 4.31 $\pm$ 0.56 Pollen 2 5 5 15 4.31 $\pm$ 0.67 Satellite 2 36 36 1254 4.84 $\pm$ 0.44 Scene 2 299 304 19178 8.61 $\pm$ 1.27 Speed Dating 2 120 500 3454 7.12 $\pm$ 1.16 Telco Customer Churn 2 19 45 831 5.52 $\pm$ 0.65 Vehicle sensIT 2 100 100 1278 4.66 $\pm$ 0.83 waveform-5000 2 40 40 306 4.62 $\pm$ 0.65